OpenStudy (anonymous):

(-6i)(-6i) a. -36i b.36 c.-36 d.36i

OpenStudy (anonymous):

c

OpenStudy (anonymous):

-36

OpenStudy (niel):

the reason why:

OpenStudy (anonymous):

i^2 = -1 -6 * -6 = 36 so -36

jhonyy9 (jhonyy9):

36i*2 because i*2 is -1 will be -36

OpenStudy (anonymous):

i had got d

OpenStudy (niel):

do the math -6x-6 is 36 i x i is i^2 which makes -1 so 36 x -1

OpenStudy (anonymous):

by definition i = sqrt of -1

OpenStudy (anonymous):

\[its -36\]

OpenStudy (anonymous):

do u get it kates?

OpenStudy (anonymous):

no. neg times a neg is a positive right?

OpenStudy (anonymous):

thats true for real numbers but not for imaginary numbers 'i' is an imaginary number.

OpenStudy (niel):

yes but the imaginary when multiplied by itself makes a third negativ e

OpenStudy (anonymous):

the number i was assumed to exxist in order to solve equations like x^2+1 = 0